OpenStudy (anonymous):

6(7m - 6) + 6 = 306 Please explain step by step It's called multi step with distribution

Nnesha (nnesha):

distribute parenthesis by 6

OpenStudy (anonymous):

Ok (42m - 36)

Nnesha (nnesha):

yes right now you can remove that parenthesis

OpenStudy (anonymous):

Ok

Nnesha (nnesha):

\[\huge\rm 42m -36 + 6 = 306\] combine like terms on left side

OpenStudy (anonymous):

Ok so you plus 6 to the 36 right?

Nnesha (nnesha):

YES but 36 is negative so -36 + 6 = ?

OpenStudy (anonymous):

So..... 42m - -30 now?

Nnesha (nnesha):

-36 + 6 = -30 so there should be one negative sign \[\huge\rm 42m - 30 = 306\] now solve for m

OpenStudy (anonymous):

Ok 42m - 30 = 306 -42 -42 = 254

Nnesha (nnesha):

nope 42 is multiplying with m so first step to move -30 to the right side so add 30 both side first then divide by 42

OpenStudy (anonymous):

Ohhhhh

OpenStudy (anonymous):

So it will be 336

Nnesha (nnesha):

final answer ??

OpenStudy (anonymous):

M= 336

Nnesha (nnesha):

nope when yo uadd 30 both side yo uget 336 but left side u still have 4m\[\huge\rm 4m = 336\]

Nnesha (nnesha):

now divide both side by 4

Nnesha (nnesha):

what you get ?

OpenStudy (anonymous):

M=84

Nnesha (nnesha):

336/4 =?

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

Thx so much

Nnesha (nnesha):

wait

Nnesha (nnesha):

hmm 6 times 7m =42 i did a mistake sorry my bad \[\huge\rm42m = 336\] divide by 42 not 4 sorry

OpenStudy (anonymous):

Yeah I had it right the first time lol

Nnesha (nnesha):

nope what's the answer now ?

OpenStudy (anonymous):

We'll kinda

OpenStudy (anonymous):

Ummm

Nnesha (nnesha):

 that's what we got after combining like terms right